A 2kg iron block is taken from a furnace where its temperature was 650 C and placed on a large block of ice at 0C. assuming that all the heat given up by the iron is used to melt the ice, how much ice is melted?
1
Expert's answer
2012-12-19T09:33:20-0500
M - mass of melted ice

m - mass of iron (2 kg)

c - specific heat of iron

L - Enthalpy of fusion of ice

Iron cold from 650 C to 0 C and give all heat to the ice.Q1=Q2;

L*M=m*c*(t2-t1)

M=m*c*(t2-t1)/L=(460*2*(650-0))/(3.3*10^5)=1.812 kg